OpenStudy (anonymous):

What are the square roots of 0=9x^2-49 explain plz

OpenStudy (anonymous):

Start by adding 49 to both sides, you get \(9x^2=49\), and then divide both sides by 9, which gives \(x^2= \frac{49}{9}\). Taking the square roots of both sides gives the two solutions \(x= \frac{7}{3}\) and \(x= -\frac{7}{3}\).

OpenStudy (anonymous):

Thanxx

OpenStudy (anonymous):

You're welcome!